OpenStudy (anonymous):

Let Y be a function of X, given by the table of values below. X: 1.25| 1.5, |1.75| 2.0 | 2.25| 2.50 | Y: 6.25| 6.63|7.03| 7.46| 7.91 | 8.40 | (a) Verify that for constant changes in X , the change in the function is proportional to the Y value, at least at the same precision level as the given data (to two decimal places). (b) Using this value, find an approximate formula for Y in terms of X . By doing +A to each X to get Y, I noticed that: The +A was 5, 5.13, 5.28, 5.46, 5.66, 5.9 So I thought the equation must be X"something" +5, thus not sure on what variable change is used.

OpenStudy (anonymous):

I've tried finding the slope of the final X and Y values - the first X and Y values and got y=mx+b as y=1.72*x+4.1 but it doesn't work for the other X values.

OpenStudy (anonymous):

The slope is change in y over change in x. Looking at the y differences, the average is roughly 0.43 so the slope is 0.43 over 0.25 is roughly 1.72 Then use that and the average of X and Y values in Y= mX +C in order to find C (looks close to 4)

OpenStudy (anonymous):

I tried that earlier, you'll see in my previous post that is what I got, I did 7.03=1.72x1.75+4, but when I do 1.72x1.75+4 I get 7.01 which is not the original y value of 7.03. And the question asks the same precision level, so I am confused on if this is good enough.

OpenStudy (anonymous):

ahh but using 4.02 works for +C, thanks for the help. I must of been punching in the wrong numbers somewhere along the line.

OpenStudy (anonymous):

Thouhg the same C value doesn't seem to work for each one, 1.72x1.25+4.02 = 6.17 when it should be 6.25. :S

OpenStudy (anonymous):

You would not expect to get perfect accuracy with a real set of figures so any equation is just going to be a best fit with some level of accuracy. In the example you gave, you are off by 0.02 which is only a small error especially when you are working to 2 decimal places

OpenStudy (anonymous):

but for X as 2 it works to make y = 7.46.

OpenStudy (anonymous):

Ahh okay, that makes sense. Cheers.

OpenStudy (anonymous):

ur welcome